(4x-1)[sup]2[/sup]=9/16
1
Expert's answer
2011-05-05T07:40:10-0400
(4x-1)2 = 9/16
4x-1 = 3/4 4x-1 = -3/4
x = 7/16 x = 1/16

(7/16, 1/16)